OpenStudy (anonymous):

A line has a rise of 6 and a slope of 1/20. What is the run?

OpenStudy (quietus):

slope = rise/run 1/20 = 6/run - cross multiply

OpenStudy (quietus):

therefore run = 6 x 20, = 120.